the external dimensions of a closed cylindrical water tank made of concrete materials consist of its radius which is 3 meters and height which is 8 meters if the thickness of the concrete wall is 1decimeter, what is the volume of the concrete walls to the nearest m3
Answer by Theo(13342)

the volume of a cylindrical prism is v = pi * r^2 * h

v is the volume
pi is the constant of 3.14.....
r is the radius
h is the height.

you are given the external dimensions of the cylinder as:

r = 3 meters
height = 8 meters.

you are also given that the thickness of the concrete walls is 1 decimeter.

you need to convert decimeters to meters or meters to decimeters.

from the conversion charts, you get 1 decimeter = 1/10 of a meter = .1 meters.

you also get 1 meter = 10 decimeters.

we'll convert everything to meters.

your outer dimensions are:

h = 8 meters
r = 3 meters.
thickness of the wall is 1 decimeter = .1 meters.

your inner dimensions are:
h = 8 meters (no change here)
r = 2.9 meters (subtract .1 meter from 3 meters to get 2.9 meters)

the volume of the external dimensions of the tank is equal to pi * 3^2 * 8 which is equal to 72 * pi cubic meters.

the volume of the internal dimensions of the tank is equal to pi * (2.9)^2 * 8 which is equal to 67.28 * pi cubic meters.

the volume of the concrete wall is there the difference which is 72 * pi minus 67.28 * pi = 4.72 * pi.

your solution is:

the volume of the concrete walls, rounded to the nearest cubic meters, is equal to 5.